birth weights are normally distributed with the mean of 3570g and a standard deviation of 495g. if a hospital wants to watch carefully the lightest 2% of infants, find the weight that seperates the bottom 2% from others.
Answer by stanbon(48568)

birth weights are normally distributed with the mean of 3570g and a standard deviation of 495g. if a hospital wants to watch carefully the lightest 2% of infants, find the weight that seperates the bottom 2% from others
----------------------
Sketch a normal curve and put a point on the horizontal axis
that has 2% of the area to the left under the curve.
-----------
The z-value of that point is -2.0537..
----------
Find the raw score (x-value) associated with that z-score, using
z= (x-u)/s
-2.0537 = (x-3570)/495
x-3570 = -2.0537*495
x = -1016.61 + 3570
x = 2553.39g
=================
Cheers,
Stan H.